From owner-hst-status@stsci.edu Fri Mar 27 09:34:50 1998 X-Sender: cwolf@mail.hst.nasa.gov To: george.albright@hq.nasa.gov, sadams@estif2.gsfc.nasa.gov, barnbaum@stsci.edu, preston.m.burch.1@gsfc.nasa.gov, jconwell@ems.jsc.nasa.gov, john.h.campbell.1@gsfc.nasa.gov, bcooke@sssp.hst.nasa.gov, carpenter@tma1.gsfc.nasa.gov, aferro@as.arizona.edu, jeffrey.m.hanley1@jsc.nasa.gov, hst-status@stsci.edu, chenning@pop500.gsfc.nasa.gov, gsfc.hst.general@newsfeed.gsfc.nasa.gov, j.keith.kalinowski.1@gsfc.nasa.gov, ken.ledbetter@hq.nasa.gov, j_lesko@swales.com, jlegg@stx.com, jleibee@hst.nasa.gov, dleckrone@hst.nasa.gov, ann.c.merwarth.1@gsfc.nasa.gov, jmccullo@ems.jsc.nasa.gov, gmorrow@hst.nasa.gov, mgw@resource-intl.com, mnolan@pop100.gsfc.nasa.gov, Jan.Owings@gsfc.nasa.gov, kpulkkinen@v2pop.hst.nasa.gov, gilbert.roth@hq.nasa.gov, rrasche@as.arizona.edu, arankin@pop400.gsfc.nasa.gov, steven.l.smith1@jsc.nasa.gov, stdesk@eso.org, rthompson@as.arizona.edu, edward.weiler@hq.nasa.gov, vweyers@pop400.gsfc.nasa.gov, howard.j.wood.1@gsfc.nasa.gov Subject: Corrected Daily Report #2089 03/28/98 ***See 2.3 for corrections to STIS status*** HUBBLE SPACE TELESCOPE DAILY REPORT #2089 PERIOD COVERED: 0000Z (UTC) 03/26/98 - 0000Z (UTC) 03/27/98 Daily Status Report as of 086/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed NIC/2 7226 (Search for Massive Jupiters) The Near Infrared Camera (Camera Two) was used to make a corographic search for massive planets around nearby, young main sequence star Rosat7. Because of the extreme youth of this object, any low-mass brown dwarf and planetary companions will still be in a higher luminosity phase and thus easily detectable. A dark calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed TWo Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed NIC/1/2/3 7245 (Galaxy Evolution in Rich Clusters) The WF/PC-2, the Near Infrared Camera (Cameras One, Two and Three) and Near Infrared Camera were used to make observations of the galactic cluster Abell-1795. The purpose is to study the morphological and spectral evolution of galaxies within this cluster environment. A dark calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.4 Completed FOC/96 7923 (FOC F/96 Turnon) The Faint Object Camera (f/96) was used to take an internal flat and a dark to ensure that the 30 day limit for the FOC is not exceeded. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed NIC/1/2/3 7328 (High Resolution IR Imaging Survey (IRIS) of the Centers of the Nearest Active Galaxies) The Near Infrared Camera (Camera Ones, Two and Three) was used to make high resolution infrared observations of the nearest active galaxy Markarian-1 in order to study its inner nuclear region. Observing in the near-IR minimizes the dust obstruction and maximizes sensitivity to the old stellar population which dominates the mass. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed STIS/CCD 7593 (CCD Spectroscopy of the Red Rectangle) The Space Telescope Imaging Spectrograph (CCD) was used to make spectrographic images of star HD44179 (a.k.a., The Red Rectangle). This star is a central star with dense, dusty disk and surrounding optical/molecular envelope. The high spatial resolution and coronographic capabilities of the STIS will be exploited to investigate the molecular and dust formation processes in the nebula. A wavelength calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Three Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.9 Completed WF/PC-2 7622 (WF/PC-2 Cycle 7 Internal Monitor) The WF/PC-2 was used to make a set of Cycle 7 internal monitoring calibration measurements. A variety of internal exposures are obtained in order to provide a monitor of the integrity of the CCD camera electronics in both bays {gain 7 and gain 15}, a test for quantum efficiency in the CCDs, and a monitor for possible buildup of contaminants on the CCD windows. A bias calibration and an internal flat calibration image were also taken. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed WF/PC-2 7385 (Low-Mass Halo Binaries) The WF/PC-2 was used to make observations of the low mass, halo, binary, {Population II}, M-type, sub-dwarf, star LHS174. This star is too faint to be resolved through ground-based speckle interferometry, thus HST represents the only means of probing separations. The observations were completed as planned, and no anomalies were reported. 1.11 Completed STIS/CCD 7803 (Daily Darks to Update Acquisition Bad Pixel Table Part III) The Space Telescope Imaging Spectrograph (CCD) was used to make daily dark calibration updates. These daily darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were completed as planned, and no anomalies were reported. 1.12 Completed NIC/2 7213 (High Resolution Imaging of "Warm" Hyperluminous Infrared Galaxies) The Near Infrared Camera (Camera Two) was used to make high resolution infrared images of the "warm" hyperluminous galaxy F15307+325 in order to expand our knowledge of their complicated relationship with the QSO phenomenon. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed Two Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.14 Completed NIC/1/2 7329 (The Nature of the Damped LyAlpha Absorbers -- A New Study of Young Galaxies) The Near Infrared Camera (Cameras One and Two) was used to make observations of the young galaxy Q1209+0919 in order to study the nature of its damped Lyman-Alpha Absorbers. The damped Lyman-Alpha absorption line systems {DLAs}, occasionally found in quasar spectra, provide one of only a few practical methods of detecting and studying significant numbers of galaxies at high redshifts. The observations were executed as scheduled, and no anomalies were noted. 1.15 Completed NIC/2 7420 (A Near-Infrared Search for Very Low Mass Companions to Stars Within 10 Parsecs of the Sun) The Near Infrared Camera (Camera Two) was used to make searching observations for very low mass companions to variable star BD+20D2465. This star is known to be within 10 parsecs of the Sun. The observations were executed as scheduled, and no anomalies were noted. 1.16 Completed STIS/CCD 7809 (Prism Sensitivity and Faint Calibration Standard Extension) The Space Telescope Imaging Spectrograph (CCD) was used to make prism sensitivity and faint calibration standard extension observations using faint white dwarf star HS2027+065. This program is the basic sensitivity measurement for the prism. It is run once a year. Sensitivity measurements are done for the prism at a central wavelength of 1200 using the 25MAMA and the 2x2 apertures. Cross-calibration via all first-order L gratings is done as well. The observations were completed as planned, and no problems were noted. 1.17 Completed WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.18 Completed WF/PC-2 6540 (Imaging Circumstellar Nebulae Around Luminous Blue Variables in the Magellanic Clouds) The WF/PC-2 was used to make observations of the curcumstellar nebula around the luminous blue variable star HD269929 which is located in one of the Magallanic Clouds. The objectives are to study the distribution of the ionized gas and dust in order to determine the morphology and physical properties of the circumstellar nebulae, by observing them in a narrow, line-free V continuum filter and in the H Alpha emission line. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 10 Successful : 10 Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 21 Successful : 21 2.3 Operations Notes: Utilizing ROP SR-1, the EDAC error counter was cleared at 085/1416Z, 085/1541Z, 085/2058Z, and 0860249Z. At 085/215626Z, while in a ZOE, the MAMA2 MCE2 high voltage was autonomously turned off when the event threshhold voltage limit was exceeded. HSTAR 6411 was written. ROP NS-14 was utilized to perform a dump of the STIS error log at 086/0248Z. The ROP has not yet been completed, and STIS is not yet recovered. As described in HSTAR 6412, the STIS instrument suspended at 085/235312Z. A STIS memory dump was executed at 086/0321Z. It is expected that STIS will be recovered later this morning after analyses are completed. Science proposals affected by the suspend will be reported in the next scheduled report. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH